OpenStudy (anonymous):

A blue bike is $14 less than a red bike. The sum of their prices is $300. How much is the red bike?

OpenStudy (anonymous):

Red Bike Cost is $ 157.

OpenStudy (anonymous):

B+R=300 B=R-14 R-14+R=300 2R=314 R=157